1. A 120cm long uniform rod is balanced at point P with a mass of 125g at one of its ends as shown. Find the mass of the rod
1
Expert's answer
2021-11-11T10:37:49-0500




"mg\cdot 0.3+F_1\cdot0.15=F_2\cdot0.45"


"mg\cdot 0.3+\frac{M}{1.2}\cdot0.3\cdot g\cdot0.15=\frac{M}{1.2}\cdot0.9\cdot g\cdot0.45"


"0.125\cdot 0.3+\frac{M}{1.2}\cdot0.3\cdot0.15=\frac{M}{1.2}\cdot0.9\cdot0.45\to"


"M=0.125\ (kg)" . Answer